The perimeter of a rectangle is 54 inches. Find the length of the rectangle if the width is 10 inches.
Answer by rm29924(97)

the formula for the perimeter of a rectangle is
perimeter=2(width+height(same thing as length))
we know the perimeter and the width so we plug it in the formula
let L=length
54=2(10+L)
54=20+2L
34=2L
17=L
check your answer:
54=2(10+17)
54=54